An aeroplane of mass 1.12 ×10^5 kg is accelerating constantly from rest along a runaway. Calculate the weight of the aeroplane. Name one of the two forces acting on the aeroplane apart from weight. The aeroplane starts from rest and after 50 seconds the aeroplane reaches a speed of 70m/s. Calculate the horizontal force on the aeroplane


1
Expert's answer
2022-05-23T10:17:59-0400

The Newton's second law says

"F=ma=m\frac{\Delta v}{\Delta t}"

"F=\rm 1.12*10^5\: kg*\frac{70\: m/s}{50\: s}=1.6*10^5\: N"